OpenStudy (rootbeer003):

What value should be added to the expression to create a perfect square? x^2 + 16x 8 16 64 4

OpenStudy (rootbeer003):

was thinkin 16

OpenStudy (rootbeer003):

wait no 4

OpenStudy (anonymous):

two steps only and no, not 16 no not that either

OpenStudy (anonymous):

what is half of 16? that is step one

TheSmartOne (thesmartone):

Half of 16 And square it

OpenStudy (rootbeer003):

8

OpenStudy (anonymous):

second step what is \(8^2\)?

OpenStudy (rootbeer003):

64

OpenStudy (anonymous):

done finished next ...

TheSmartOne (thesmartone):

Bingo

OpenStudy (rootbeer003):

hold up

OpenStudy (anonymous):

holding...

OpenStudy (anonymous):

\[2x^2-2x-1=0\] solve via \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=2,b=-2,c=-1\] or cheat

OpenStudy (rootbeer003):

so i right?

OpenStudy (anonymous):

yes

OpenStudy (rootbeer003):

c?

OpenStudy (anonymous):

\[b^2-4ac\] with \(a=1,b=-3,c=5\) i.e. \[(-3)^2-4\times 1\times 5\]

OpenStudy (anonymous):

oh yea, C

OpenStudy (anonymous):

what is \(b^2-4ac\) here?

OpenStudy (rootbeer003):

wat

OpenStudy (anonymous):

let me know when you get \(36-12=24\)

OpenStudy (anonymous):

important part is that it is a positive number that means there are two real solutions

OpenStudy (anonymous):

why the stupid restrictions? just graph it here

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=-2x^2%2Bx-3

OpenStudy (anonymous):

only question is, can you read the graph to answer the question is the vertex above or below the x axis?

OpenStudy (anonymous):

i need a drink, brb

OpenStudy (rootbeer003):

above?

OpenStudy (anonymous):

does it look like it is above to you !!??

OpenStudy (anonymous):